Shift report 20030227 00:00-08:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Thu Feb 27 2003 - 07:54:50 EST

  • Next message: Hironori Ito: "c4 and TOFW2 detectors on the web"
    ------------------------------------------------------------
    Date: Thu Feb 27 07:54:50 2003, Shift: 00:00-08:00
    Supervisor: MM, with: 
    Log book#14 pages 37 - 37
    
    Shift Summary:
    ------------------------------------------------------------
    1.RHIC cooling failure. There has been a cooling problem in one of the RHIC magnets, that has lasted the entire night. Repairs are still beeing done. No beam during shift.
    
    2.Only pedestal and DVM run was taken. Exept tpc pedestal run that I could not do. (eventrate very very low)
    
    3.No detector problems observed. 
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Thu Feb 27 2003 - 07:56:01 EST